OpenStudy (anonymous):

find the equation of the tangent to the circle x^2+y^2=25 drawn through the point(13,0).

OpenStudy (anonymous):

\[\frac{dy}{dx}=-\frac{x}{y}\] so the slope between the point \((x,y)\) on the circle and \((0,13)\) must satisfy the equation \[\frac{y-13}{x}=-\frac{x}{y}\] and since it is on the circle they must also satisfy the equation \(x^2+y^2=25\)

OpenStudy (anonymous):

solve the two equations to get your answer

OpenStudy (anonymous):

well i got answer 12y=+ and -5(x-13) is it true?

OpenStudy (anonymous):

oh damn it is \((13,0)\) not \((0,13)\) ok idea is still the same but equation is wrong \[\frac{y}{x-13}=-\frac{x}{y}\]

OpenStudy (anonymous):

well i still didn't get it what u r trying to say

OpenStudy (anonymous):

maybe i am mistaken. you want the equation of the line tangent to the circle i took the derivative of \(x^2+y^2=25\) with respect to \(x\) to find the equation for the slope. are you using a different method?

OpenStudy (anonymous):

yes i guess so

OpenStudy (anonymous):

i just checked and your answer looks correct

OpenStudy (anonymous):

just for my curiosity, what method are you using?

OpenStudy (anonymous):

whatever you did, it worked nice picture here http://www.wolframalpha.com/input/?i=12y%3D5%28x-13%29%2C+x^2%2By^2%3D25%2C12y%3D-5%28x-13%29

OpenStudy (anonymous):

well the center of the circle is(0,0) and the radius is 5 so the equation throught the point (13,00 is simply y=m9x-13) and where the m is the slope and the lenght of the perpendicular drawn from the centre on the line == and - 0-0=13m/suare root (1+m^2) and therefore m= +and - 5/12 and by substituting i got the answer